OpenStudy (anonymous):

Suppose that f(x) = frac{1}{x}. Show that there is no real number c in [-1,2] such that frac{f'(c) = f(2)-f(-1)}{2-(-1)}.

OpenStudy (anonymous):

\[f(x) = \frac{1}{x}\] Show that there is no real number c in [-1,2] such that \[f'(c) = \frac{f(2)-f(-1)}{2-(-1)}\] Why does this not contradict the mean value theorem? f(x) is not continuous on [-2,2].

OpenStudy (anonymous):

I don't really understand how to approach this problem.

OpenStudy (anonymous):

work strictly with numbers on the right the derivative on the left is \(-\frac{1}{x^2}\)

OpenStudy (anonymous):

I solved for c and got \[c = \pm \sqrt{2}\] I'm not sure how this shows there are no real roots in [-1,2] though, as it seems there is at least one! Right?

OpenStudy (anonymous):

i hope it is clear what i mean the right hand side of the equal sign is a number what did you get?

OpenStudy (anonymous):

Ah never mind, I screwed up the arithmetic. I think it ends up as \[x^{2} = -2\] so there are no real roots, only complex.

OpenStudy (anonymous):

yes, and so no solution

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

and your answer above is correct, it is not continuous on that interval